Solve Problems Using Linear Systems
14. A metal alloy is 25% copper. Another metal alloy is 50% copper. How much of each should be used to make 500 g of an alloy that is 45% copper?
(The answer is: 100 g of metal alloy that is 25% copper, 400 g of metal alloy that is 50% copper, but i dont know how to solve it?)

Answer by checkley77(12844)

.50X+.25(500-X)=.45*500
.50X+125-.25X=225
.25X=225-125
.25X=100
X=100/.25
X=400 G OF 50% COPPER IS USED.
500-400=100 G OF 25% COPPER IS USED.
PROOF:
.50*400+.25*100=.45*500
200+25=225
225=225